OpenStudy (anonymous):

inverse laplace transform F(s)=2s+12/s^2+2s+5

OpenStudy (anonymous):

\[F(s)=\frac{2s+12}{s^2+2s+5}\] I would first complete the square in the denominator, then split up into two fractions: \[s^2+2s+5\\ s^2+2s+1+4\\ (s+1)^2+4\] \[F(s)=\frac{2s}{(s+1)^2+4}+\frac{12}{(s+1)^2+4}\] If you have a table of inverse transforms near you, it helps to rewrite the above as \[F(s)=\frac{2s+2-2}{(s+1)^2+4}+\frac{12}{(s+1)^2+4}\\ F(s)=2\left(\frac{s+1}{(s+1)^2+4}\right)+\frac{10}{(s+1)^2+4}\\ F(s)=2\left(\frac{s+1}{(s+1)^2+4}\right)+5\left(\frac{2}{(s+1)^2+4}\right)\] Does that help?

OpenStudy (anonymous):

@raprap, if you don't have a table of inverse transforms, you should know that, if \(\mathcal{L}\{f(t)\}=F(s)\), then \[\mathcal{L}^{-1}\left\{F(s-a)\right\}=e^{at}f(t)\\ \mathcal{L}^{-1}\left\{\frac{s}{s^2+b^2}\right\}=\cos t\\ \mathcal{L}^{-1}\left\{\frac{b}{s^2+b^2}\right\}=\sin t\]